if it takes 24.8 ml of 1.20N acid to neutralize 3.25 g of a base, what is the equivalent weight of the base

1
Expert's answer
2022-05-19T05:19:02-0400

1. Using the formula CN = nequ / V, we obtain the formula for finding nekv: nequ = CN * V. where,

CN – is the normal concentration,

nequ – is the equivalent amount of the substance,

V – is the volume.

nequ = CN * V = 1,2 * 0.0248 = 0.02976

hence the equivalent amount of reacting acid is 0.02976. 

2. substances react in equal equivalent amounts with each other. therefore the equivalent amount of reacting base is also 0.02976.

3. Using the Mequ = m / nequ formula, we find the equivalent mass of the acid. To do this, we divide the mass of the acid by its equivalent amount:

Mequ = m / nequ = 3,25 / 0.02976 = 109.2

ANSWER: 109.2 gr/equivalent